# Prompt-based Depth Pruning of Large Language Models

### v2

@@ 1 Introduction
Table 1: A high-level comparison of the proposed prompt-based depth pruning framework with related approaches: Static depth pruning and dynamic token routing.
@@ 2.2 Dynamic Token Routing
2.3 Side Note: Width Pruning Techniques
Width pruning, i.e., pruning out columns and rows of the weight matrices in a structured manner, has also been popularly used for LLM compression. Recent examples include FLAP (An et al., 2024) and SliceGPT (Ashkboos et al., 2024) for the category of static width pruning, and Déjà Vu (Liu et al., 2023) for the dynamic width pruning.
As these methods tend to alter the size of the weight matrices, which has been highly customized for hardware considerations, they often lead to further challenges in deployment (Song et al., 2024; Kim et al., 2024). In this work, we thus mainly focus on comparison with depth pruning, but also compare with width pruning in terms of the accuracy.
@@ 3 A Motivating Observation
Result. Figure 2 illustrates our findings. We observe that pruning out block 29 instead of block 30 has a two-sided impact: On BoolQ, the change makes a dramatic drop in accuracy (62.2% $\to$ 38.0%, 62.5% $\to$ 37.9%). However, on PIQA and WinoGrande, we observe a slight accuracy boost. This phenomenon suggests that the block 30 may contain more knowledge relevant to answering BoolQ questions, while 29 may be more knowledgeable about PIQA and WinoGrande. This observation highlights the need to consider task variability during the selection of the omission set. To formally address such need, this paper considers an inference of task information from the prompt.
@@ 6 Experiment
PuDDing (Ours) Depth 7 ($>$21%) 41.47 67.09 62.02 62.92 73.94 64.16 61.93 (+2.95)
@@ 6.1 Experimental Setup
Dataset: Evaluation. We evaluate on the test splits of six zero-shot commonsense reasoning tasks: ARC-Challenge and ARC-Easy (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), and BoolQ (Clark et al., 2019).
Dataset: Calibration for Baselines. For the baseline algorithms, we have used the calibration data designated in the original papers. For SLEB, FLAP, and SliceGPT, we have used the WikiText-2 (Merity et al., 2022). For Shortened LLaMA, we have used the BookCorpus (Zhu, 2015).
@@
Pruned Blocks
OpenBookQA MathQA MMLU
Dense 0 44.60 39.53 52.71
@@
PuDDing 7 ($>$21%) 36.40 27.20 33.63
: averaged over five subtasks
@@ 6.4 More Complicated Tasks
In Table 5, we compare the performance of various depth pruning algorithms on more complicated tasks, including OpenBookQA (Mihaylov et al., 2018), MathQA (Amini et al., 2019), and MMLU (Hendrycks et al., 2021); for MMLU, we take an average on five key tasks: High-school math, physics, medicine, machine learning, and philosophy. From the results, we observe that PuDDing continues to perform better than the baselines, even though these tasks have not been observed during the training of the router.
@@ 7.1 Wall-clock Speedup
Inference. Table 6 presents the average wall-clock inference time comparison between the dense and PuDDing-pruned version of the LLaMA 3.1 8B, evaluated on NVIDIA A100 and RTX 6000 Ada. For PuDDing, we have pruned seven layers (21.88% sparsity). We observe that PuDDing provides a consistent 1.19-1.23$\times$ speedup during the pre-fill stage, and 1.22-1.25$\times$ speedup including the generation stage. The total routing time takes up to 4-8ms, which can be deemed negligible comparing with the overall latency.
@@ 8 Conclusion
9 Impact Statement

### v3

@@ Abstract
Project Page: jwee01.github.io/PuDDing
Code: github.com/tada0347/PuDDing
@@ 1 Introduction
Table 1: A high-level comparison of the proposed prompt-based depth pruning framework with related depth pruning approaches: Static depth pruning and dynamic token routing.
@@ 2.2 Dynamic Token Routing
2.3 Contextual Sparsity
Our work is most closely related to the idea of contextual sparsity, where a lightweight router selects an input-dependent subnetwork at inference time without updating the base weights. In the context of width pruning, prior works—Deja Vu (Liu et al., 2023), ShadowLLM (Akhauri et al., 2024), Sirius (Zhou et al., 2024), and CATS (Lee et al., 2024)—have demonstrated that context-aware routing can be done with minimal or no degradation in task performance. PuDDing extends this paradigm to depth pruning for the first time: instead of skipping neurons or channels, our router decides which entire transformer blocks to omit. This preserves the original matrix shapes and avoids hardware mismatches often caused by width pruning.
@@ 3 A Motivating Observation
Result. Figure 2 illustrates our findings. We observe that pruning out block 29 instead of block 30 has a two-sided impact: On BoolQ, the change makes a dramatic drop in accuracy (62.2% $\to$ 38.0%, 62.5% $\to$ 37.9%). However, on PIQA and WinoGrande, we observe a slight accuracy boost. This phenomenon suggests that the block 29 may contain more knowledge relevant to answering BoolQ questions, while 30 may be more knowledgeable about PIQA and WinoGrande. This observation highlights the need to consider task variability during the selection of the omission set. To formally address such need, this paper considers an inference of task information from the prompt.
@@ 6 Experiment
PuDDing (Ours) Depth 7 ($>$21%) 41.47 67.09 62.02 62.92 73.94 64.16 61.93 (+4.69)
@@ 6.1 Experimental Setup
Dataset: Evaluation. We evaluate on the test splits of six zero-shot commonsense reasoning tasks: ARC-Challenge and ARC-Easy (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020), and BoolQ (Clark et al., 2019).
Dataset: Calibration for Baselines. For the baseline algorithms, we have used the calibration data designated in the original papers. For SLEB, FLAP, and SliceGPT, we have used the WikiText-2 (Merity et al., 2022). For Shortened LLaMA, we have used the BookCorpus (Zhu et al., 2015).
@@
Pruned
Blocks
OpenBookQA MathQA MMLU PubMedQA SciQ
Dense 0 44.60 39.53 63.49 75.80 96.00
@@
PuDDing 7 36.40 27.20 39.00 60.00 92.70
@@ 6.4 More Complicated Tasks
In Table 5, we compare the performance of various depth pruning algorithms on more complicated tasks, including OpenBookQA (Mihaylov et al., 2018), MathQA (Amini et al., 2019), and MMLU (Hendrycks et al., 2021). From the results, we observe that PuDDing continues to perform better than the baselines, even though these tasks have not been observed during the training of the router.
@@ 7 Analysis
Table 7: Wall clock inference speed of the PuDDing-compressed LLaMA-3.1 8B evaluated on edge devices (M3 pro, Apple).
M3 Pro (Apple) Pre-fill (TTFT) Pre-fill + Generation
Prompt Length 128 256 512 128 128 128
Gen. Length 1 1 1 128 256 512
Dense 0.177s 0.300s 0.480s 7.890s 15.970s 32.520s
PuDDing 0.138s 0.235s 0.376s 6.174s 12.497s 25.447s
→ Router 0.009s 0.016s 0.029s 0.009s 0.009s 0.009s
Speed Up 1.20× 1.20× 1.19× 1.28× 1.28× 1.28×
@@ 7.1 Wall-clock Speedup
Inference. Table 6 presents the average wall-clock inference time comparison between the dense and PuDDing-pruned version of the LLaMA 3.1 8B, evaluated on NVIDIA A100 and RTX 6000 Ada. For PuDDing, we have pruned seven layers (21.88% sparsity). We observe that PuDDing provides a consistent 1.19-1.23$\times$ speedup during the pre-fill stage, and 1.22-1.25$\times$ speedup including the generation stage. The total routing time takes up to 4-8ms, which can be deemed negligible comparing with the overall latency. Also, Table 7 presents results on edge devices (e.g., Apple M3 Pro), showing consistent speedup. This outcome shows that the proposed method is well-suited for both server-like and edge-like hardwares.
@@ 8 Conclusion
Acknowledgements
This work has been supported by National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) (Nos. RS-2023-00213710, RS-2024-00453301).
Impact Statement
@@ B.3 Using MSE Loss for Training
B.4 Training the Router with Multi-Domain Calibration Datasets
Table 13: Zero-shot accuracy comparison of dense, PuDDing, and PuDDing-MultiDomain using diverse calibration datasets.
Method Average Acc. (%) Per-task Accuracies
Arc-C Arc-E BoolQ HellaSwag PIQA WinoGrande MathQA PubMedQA SciQ
Dense 73.42 53.50 81.52 82.20 78.81 79.98 73.40 39.53 75.80 96.00
PuDDing 61.28 41.47 67.09 62.02 62.92 73.94 64.16 27.27 60.00 92.70
PuDDing-MultiDomain 62.37 41.38 67.26 67.37 63.68 73.07 64.56 29.58 62.00 92.40
In Section 6, we construct omission sets using common-sense reasoning datasets that are widely used to evaluate language models’ reasoning ability—ARC, PIQA, HellaSwag, and WinoGrande—to reflect popular usage scenarios. For the new variant, PuDDing-MultiDomain, we conduct an additional experiment by incorporating calibration datasets from diverse domains: MathQA (mathematics), PubMedQA (biomedical), and SciQ (science). To maintain consistency in sample size with the original setting, we include only a subset of the original common-sense datasets—specifically, ARC-Easy and WinoGrande—in this experiment. All other experimental settings follow the same setup described in Section 6.1.
We observe that PuDDing-MultiDomain achieves slightly improved average accuracy compared to the original PuDDing, with noticeable gains in the newly introduced datasets (e.g., +2.31%p in MathQA, +2.00%p in PubMedQA). This indicates that the router can benefit from more diverse calibration dataset, especially when deployed in tasks requiring domain-specific knowledge. Overall, our method demonstrates consistent performance across both in-domain and out-of-domain tasks, and can be further customized to specific user applications by adjusting the coverage of the omission set pool.
B.5 Quantized PuDDing
Table 14: combined with other compression techniques
LLaMA-3.1-8B Average Arc-C Arc-E BoolQ HellaSwag PIQA WinoGrande
Dense 74.90 53.50 81.52 82.20 78.81 79.98 73.40
SLEB 57.24 34.90 66.25 49.11 61.60 74.37 57.22
Shortened LLaMA 55.77 34.30 65.15 44.52 60.55 73.67 56.43
PuDDing 61.93 41.47 67.09 62.02 62.92 73.94 64.16
PuDDing + W8A16 (AWQ) 61.68 41.30 67.00 61.50 62.95 73.72 63.61
PuDDing + W4A16 (AWQ) 58.58 37.37 61.45 60.64 57.55 71.71 62.75
In Table 14 presents the performance of the proposed method when combined with a representative compression technique such as quantization (e.g. Activation-aware Weight Quantization (AWQ) (Lin et al., 2024)). Interestingly, we observe that compressing the weights to 8-bit results in no performance degradation. Although 4-bit quantization introduces some degradation, the performance still surpasses that of static pruning methods applied with BF16 precision.